OpenStudy (anonymous):

What is the solution of log3 x − 5 16 = 2? x = −4 x = 4 x = −3 x = 3

OpenStudy (anonymous):

@ganeshie8

OpenStudy (dan815):

should that say 5^16 or something

ganeshie8 (ganeshie8):

yeah looks like it got messed up while copy past

ganeshie8 (ganeshie8):

\(\large \log_{3x-5} ~16 = 2\) like this @gabylovesu ?

OpenStudy (dan815):

ahh that makes sense

OpenStudy (anonymous):

I guess.

ganeshie8 (ganeshie8):

start by changing it to exponent form, use below : \(\large \log_a b = c \) can be written in exponent form as \(\large b = a^c\)

ganeshie8 (ganeshie8):

\(\large \large \log_{3x-5} ~16 = 2\) can be written as \(\large 16 = (3x-5)^2\)

ganeshie8 (ganeshie8):

Notice that 16 is a special number - any guess why is it special ?

OpenStudy (dan815):

you there gaby?

OpenStudy (anonymous):

Yeah I am it is just that my computer is acting up.

OpenStudy (anonymous):

Um I am not sure why the 16 is a special number. @ganeshie8

ganeshie8 (ganeshie8):

Okay, 16 is a perfect square : 16 = 4^2

ganeshie8 (ganeshie8):

\(\large 4^2 = (3x-5)^2\)

ganeshie8 (ganeshie8):

since the exponents are equal, we can compare the bases : \(\large 4 = 3x - 5\) solve \(\large x\)